Night Report 2014-01-23

Observers: Giorgos Leloudas, Mattias Ergon

Support: Jesper Sollerman, Francesco Taddia, Joel Johansson

TAT on duty: Annalisa de Cia, Maria Teresa Botticella

--------------------------------------------------------------

WEATHER CONDITIONS:

start: wind 1.5 m/s; temp. 13.7 C; humid. 63%; Some clouds but looks like improving.

middle: wind 0.3 m/s; temp. 13.2 C; humid. 44%; Clouds have cleared out (except perhaps towards the horizon)

end: wind 4.8 m/s; temp. 11.2 C; humid. 76%; Clear.

--------------------------------------------------------------

CALIBRATIONS

Afternoon:

EFOSC bias, arcs

Twilight:

Morning : grz flats

--------------------------------------------------------------

OBSERVATIONS

Target Type UT Filter/Grism Slit Seeing Comments

L745a Spec std 00:42 g11, g16, g13 1, 1.5 1.1

SN 2013ej Follow-up 01:00 g11 1 1.0 Airmass: 2.33 @start. Almost reached pointing limit.

Image analysis. Here we also lost time with trying to place both objects in the slit.

LSQ13cuw Follow-up 02:15 g13 1 1.1 Slit through SN & nearby galaxy to determine host z

LSQ13ddu Follow-up 03:50 gr 0.9 We will do the iz filters another day.

OGLE-2013-SN-016 Follow-up 04:52 Vri 1.1 ERROR : the r-band was actually UNFILTERED !!!!!!

L745a Spec std 05:38 g11, g16, g13 1, 1.5 1.1

LSQ14in Class 06:02 g13 1 1.1 Target has moved by ~20 arcsec. So not a SN. Aborted.

Image analysis

LSQ14ie Class 06:20 g13 1 1.0 SN Ia @ -8d, z = 0.079

LSQ14gs Class 06:56 g13 1 1.3 SN Ic @ max. Good match to SN 2004aw, z~0.045

LSQ14hq Class 07:17 g13 1 1.2 SN Ia @ +5d, z= 0.078

LSQ14ge Class 07:40 g13 1 1.2 Object not here ...

LSQ14gf Class 07:45 g13 1 1.2 Object not here ... Annoying ...

MASTERJ110457.1+300122.0 Class 07:59 g13 1 1.2 SN Ia @ max, z= 0.03

LSQ14hm Class 08:12 g13 1 1.2 SN Ia @ -1d, z = 0.06

LTT3864 Spec std 08:33 g11, g16, g13 1, 1.5 1.1